OpenStudy (anonymous):

x^2+18x+?=?^2 what numbers makes this true ?

OpenStudy (anonymous):

did you copy the correct prob here?

OpenStudy (anonymous):

Yes , I have to find what number fits in thee blanks . I dont get it at all .

OpenStudy (anonymous):

x^2+18x+?=?^2 try completing the squares x^2+18x+(18/2)^2=y^2+(18/2)^2 (x+9)^2=y^2+81

OpenStudy (anonymous):

Let me see if I can first write the correct form of the problem: \[x^2+18x+a=a^2\] which implies: \[x^2+18x+a-a^2=0\] Now, we use the quadratic formula: \[x=(1/2a)(-b+or-\sqrt{b^2-4ac})\] Set a=1, b=18, and c= a-a^2=a(1-a) Using the these values for the quadratic formula from above we see that x must be one of two values: 1.\[(1/(2a(1-a))(-18+\sqrt{324-4a(1-a})\] or 2.\[(1/(2a(1-a))(-18-\sqrt{324-4a(1-a})\] Plug these values in for the "x" variable in the original quadratic equation and you will obtain the criteria that a must satisfy in order for the statement to be true.

OpenStudy (anonymous):

*sorry, at the end I said quadratic equation and it could easily be confused with the quadratic formula. Plug the two values into the original polynomial